# Unravelling the emergence of quantum jumps in a monitored qubit

### v1

@@ Abstract
Quantum jumps, the collapse of a quantum system upon measurement, are among the most striking consequences of observation in quantum mechanics. While recent experiments have revealed the continuous nature of individual jumps [21], the crossover from coherent dynamics to measurement-dominated behaviour has remained elusive. Here, we tune the measurement strength of a continuously monitored superconducting qubit, and observe that quantum jumps emerge not through a gradual crossover, but via a cascade of three distinct dynamical transitions [38]. The first transition manifests as an exceptional point [28] where coherent oscillations abruptly cease, giving way to jumps towards a stable eigenstate. The second transition marks the onset of dynamical state freezing, where the qubit’s dwell time near the eigenstate diverges. A third threshold signals entry into the quantum Zeno regime, where stronger measurement paradoxically suppresses relaxation [12]. Strikingly, we find that decoherence does not blur these transitions but rather fundamentally restructures the dynamical phase diagram, notably inverting their order. These results map measurement-induced transitions in a monitored qubit, revealing that the interplay between coherent driving, measurement, and decoherence gives rise to a hierarchy of distinct dynamical phases.
A central question in quantum mechanics is how measurement shapes the behaviour of quantum systems. While unobserved systems evolve deterministically under the Schrödinger equation, measurement introduces stochastic evolution [6] through outcome-dependent backaction [14, 41, 27]. This tension is particularly striking in a continuously monitored two-level system. Consider a qubit with ground state $\left|0\right\rangle$ and excited state $\left|1\right\rangle$, driven resonantly at a Rabi rate $\Omega_{\mathrm{S}}$ and monitored by a detector that clicks at a rate $\alpha$ when the system is in $\left|0\right\rangle$ (Fig. 1). Quantum trajectory theory [5, 25, 43] describes how stochastic detector clicks project the system to $\left|0\right\rangle$, while between clicks, the wavefunction evolves deterministically under an effective non-Hermitian Hamiltonian (Supplementary Information Section 4[1, 3, 28]
@@
Here, the dimensionless measurement strength $\lambda\equiv\alpha/(2\Omega_{\mathrm{S}})$ quantifies the competition between coherent driving and no-click backaction [32].
The behaviour at the extremes of this parameter is well established. For weak measurements ($\lambda\ll 1$), coherent driving dominates, resulting in Rabi oscillations (Fig. 1). In contrast, for strong measurements ($\lambda\gg 1$), measurement backaction dominates: the sustained absence of clicks drives the qubit towards $\left|1\right\rangle$, producing a “continuous quantum jump” [21].
@@
Figure 2: Overview of dynamical transitions in the ideal monitored qubit model. As the dimensionless measurement strength $\lambda\equiv\alpha/(2\Omega_{\mathrm{S}})$ increases, the qubit dynamics evolve through four distinct regimes (a–d), separated by three sharp transitions at $\lambda^{c}_{1-3}=\{1,\frac{2}{\sqrt{3}},2\}$. Each row displays simulated dynamics in the different regimes, showing (from top to bottom) a representative binary click record, the conditional excited state probability $P_{\left|1\right\rangle}^{\mathrm{cond}}$ of the corresponding trajectory, and the ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$. The Bloch spheres show the average qubit dwell time per unit angle $\tau_{\theta}(\theta)$ (orange curves), plotted as radial height above the unit circle (a.u.), alongside the time evolution of the ensemble-averaged state initialised at $\left|0\right\rangle$ (teal curves). a, For $\lambda<\lambda^{\mathrm{c}}_{1}$, trajectories exhibit periodic Rabi oscillations, intermittently reset to $\left|0\right\rangle$ by detector clicks, enabling the qubit to access the full range of polar angles. b, The first transition at $\lambda^{\mathrm{c}}_{1}$ marks the onset of smooth, deterministic no-click evolution towards a stable eigenstate at polar angle $\theta_{+}$ (dotted grey lines), creating a forbidden region on the Bloch sphere (grey shading). Detector clicks typically interrupt this no-click evolution before the qubit reaches the eigenstate. c, Beyond the second transition ($\lambda>\lambda^{\mathrm{c}}_{2}$), the qubit frequently reaches and freezes near the eigenstate before a click occurs, resulting in a divergent dwell time at $\theta_{+}$. d, At the third transition point ($\lambda=\lambda^{\mathrm{c}}_{3}$), the ensemble-averaged dynamics shift from oscillatory behaviour to overdamped decay with increasingly slow relaxation, marking the onset of quantum Zeno behaviour. In realistic experimental conditions, decoherence modifies both the locations and the ordering of these transitions.
@@
However, the crossover between these two regimes remains experimentally unexplored. A recent theoretical study [38] predicted that this crossover occurs through a cascade of sharp transitions, each unveiling qualitatively distinct dynamics and statistics of quantum trajectories (Fig. 2).
@@
Figure 3: Observation of the transition to the quantum jump regime. Below the first transition (bottom Bloch sphere), coherent driving (blue arrows) dominates measurement-induced backaction (red arrows) across the entire Bloch sphere, resulting in oscillations. Above it (top Bloch sphere), measurement backaction precisely balances the coherent drive at a specific angle, causing oscillations to halt and giving rise to quantum jump dynamics. a, Tomographic data showing the excited state population $P_{\left|1\right\rangle}^{\mathrm{cond}}$ as a function of the no-click sequence duration and the measurement strength $\lambda$, revealing the shift from oscillatory to jump-like behaviour. The transition is observed at $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$ (dashed red line). Each data point is the average of $1500$ experimental runs. b, One-dimensional cuts below (bottom panel, $\lambda=0.5$) and above (top panel, $\lambda=2.2$) the first transition at $\lambda$ values indicated by the dotted lines in a. The grey lines show the simulated no-click dynamics with independently-measured parameters. Shot-noise error bars are smaller than the marker size. c, Histograms of no-click durations further confirm this transition. We initialise the qubit in $\left|1\right\rangle$ (in contrast to $\left|0\right\rangle$ in a, b) to suppress the effect of errors in long no-click intervals. The grey solid lines show fits using a sum of three complex exponentials (Methods). Throughout this figure, the single-measurement integration time is $T_{\text{int}}=320\,$ns and the coherent drive strength is set at $\Omega_{\mathrm{S}}/2\pi=100\,$kHz. The tomographic measurement is preceded by a measurement to verify that the system is in the qubit manifold.
@@
In this work, we provide the first experimental observation of this transition cascade. We realise this using a superconducting circuit consisting of a system qubit coupled to an ancillary detector qubit, forming a V‑type three‑level system similar to ref. [21]. We control the measurement strength $\lambda$ by varying the amplitude of a drive resonant with the detector qubit transition (see Methods). By tuning $\lambda$, we uncover three dynamical transitions marking the emergence of measurement-dominated behaviour (Fig. 2). At an exceptional point of the non-Hermitian Hamiltonian, coherent oscillations abruptly cease and trajectories begin evolving towards a stable state—the hallmark of continuous quantum jumps [21, 17]. Beyond a second threshold, the dwell time near this state diverges: trajectories “freeze” in its vicinity. A third transition signals entry into the quantum Zeno regime, where stronger measurement counterintuitively suppresses relaxation [31, 24, 11, 12].
@@
Our analysis combines binary detector click records with conditional quantum state tomography. This allows us to reconstruct the qubit’s dynamics conditioned on specific measurement outcomes, revealing dynamical features that would otherwise be washed out in ensemble averages. Using these methods, we find that decoherence does not merely shift the critical points predicted by the idealised theory, but fundamentally alters the transitions’ character and ordering. Nevertheless, the transitions themselves remain robust features of the dynamics. These results elucidate the emergence of continuous quantum jumps and provide a map of dynamical phases in a monitored qubit.
@@
The first transition, heralding the onset of continuous quantum jumps, arises from competition between two coherent processes acting during no-click evolution (equation (1)): the Rabi drive attempts to rotate the qubit around the Bloch sphere, while the no-click measurement backaction pulls it towards the excited state $\left|1\right\rangle$. Below a critical measurement strength $\lambda<\lambda^{\mathrm{c}}_{1}$, Rabi driving dominates at all polar angles $\theta$ on the Bloch sphere, allowing the qubit to complete full periodic rotations. As $\lambda$ approaches $\lambda^{\mathrm{c}}_{1}$, the oscillations become increasingly asymmetric, reflecting the growing impact of measurement backaction (Fig. 2a). The critical point $\lambda^{\mathrm{c}}_{1}=1$ represents the threshold at which measurement backaction first becomes strong enough to halt the qubit’s rotation. For any $\lambda\geq\lambda^{\mathrm{c}}_{1}$, backaction precisely balances coherent driving at an angle $\theta_{+}(\lambda)=2\arctan\!\left(\sqrt{\lambda^{2}-1}-\lambda\right)$, creating a stable fixed point [^1] A second fixed point appears at $\theta_{-}\equiv-\pi-\theta_{+}$ [38]. However, this fixed point is unstable, so the qubit reaches it only if specifically initialised in that state.  [33]. Trajectories now execute continuous, deterministic quantum jumps from $\left|0\right\rangle$ towards the eigenstate of $\hat{H}_{\text{eff}}$, $\left|\psi_{+}\right\rangle\equiv\sin\!\left(\frac{\theta_{+}}{2}\right)\left|0\right\rangle+\cos\!\left(\frac{\theta_{+}}{2}\right)\left|1\right\rangle$, until a stochastic detector click resets the qubit to $\left|0\right\rangle$. The fixed point at $\left|\psi_{+}\right\rangle$ creates a forbidden region $(-\pi,\theta_{+}(\lambda)]$ on the Bloch sphere, which eventually encompasses the entire left hemisphere as measurement strength increases [38] (see the Bloch spheres in Figs. 2b–d).
@@
This transition coincides with an exceptional point — a degeneracy in the spectrum of the non-Hermitian no-click Hamiltonian $\hat{H}_{\text{eff}}$ where eigenvalues coalesce. Since the Hamiltonian is parity-time ($\mathcal{PT}$) symmetric, the exceptional point separates unbroken and broken $\mathcal{PT}$-symmetric phases [3]. Beyond this point, the eigenvalue difference becomes purely imaginary, causing the oscillation frequency to vanish. Similar exceptional-point transitions have been observed in other systems [28, 44, 7, 8, 40, 15].
@@
To observe this transition experimentally, we reconstruct the qubit’s evolution conditioned on the absence of detector clicks. We perform qubit tomography after detecting in real time a no-click sequence of the desired duration. Repeating this procedure across different no-click durations and measurement strengths reveals a clear shift from oscillatory to quantum jump dynamics (Figs. 3a,b). The measurements are in excellent agreement with numerical simulations of a model that incorporates decoherence (Supplementary Information Section 8). As expected, the Rabi oscillations become increasingly distorted as $\lambda$ approaches the critical value.
@@
The distribution of no-click durations independently confirms this transition. Initialising the system in $\left|1\right\rangle$, we record the waiting time to the first detector click. Since the instantaneous click rate varies as $r(\theta)\equiv\alpha\sin^{2}\!\left(\frac{\theta(t)}{2}\right)$, Rabi oscillations imprint a periodic modulation on this distribution. The data (Fig. 3c) confirm this expectation: oscillatory patterns at weak measurements give way to monotonic decay as $\lambda$ increases. From the vanishing of oscillations, we determine that the transition to quantum-jump dynamics occurs at $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$ (Methods and Supplementary Information Section 7).
@@
A defining feature of strongly measured quantum systems is dynamical freezing: the qubit should not merely approach the eigenstate $\left|\psi_{+}\right\rangle$ but remain trapped in its vicinity for extended periods. To probe this behaviour, we examine the dwell time per unit angle $\tau_{\theta}(\theta)$, defined such that $\tau_{\theta}(\theta)\Delta\theta$ gives the mean time spent in the angular interval $\theta,\theta+\Delta\theta)$. Near the fixed point at $\theta_{+}$, theory predicts a power-law scaling $\tau_{\theta}(\theta)\sim\left[\theta-\theta_{+}(\lambda)\right]^{\xi(\lambda)}$ [38], where the sign of the critical exponent $\xi$ reflects a competition between the click rate and the rate at which trajectories approach $\theta_{+}$. When clicks dominate ($\xi>0$), the dwell time vanishes at the fixed point (Fig. 2b). When the approach rate dominates ($\xi<0$), trajectories accumulate faster than clicks can reset them, and the dwell time diverges (Fig. 2c). This sign change defines a second critical threshold at $\lambda^{\mathrm{c}}_{2}=\frac{2}{\sqrt{3}}\approx 1.15$ in the ideal model (Supplementary Information Section 4), marking the onset of dynamical state freezing.
@@
We reconstruct $\tau_{\theta}(\theta)$ experimentally by initialising the qubit in $\left|1\right\rangle$, recording the angular distribution $\rho(\theta)$ of first-click events, and normalising by the instantaneous click rate $r(\theta)$ (Methods and Supplementary Information Section 4). Fig. 4a shows the resulting dwell-time distributions across a range of measurement strengths. For weak measurement strengths ($\lambda\ll 1$), the qubit completes full Rabi oscillations and $\tau_{\theta}(\theta)$ is approximately uniform. As $\lambda$ increases, a forbidden angular region emerges. Intriguingly, this occurs at $\lambda\approx 0.57$, well before oscillations cease at $\lambda^{\mathrm{obs}}_{1}=0.99$. This decoupling of the two signatures of the first transition is absent in the ideal model, where both appear simultaneously at $\lambda^{\mathrm{c}}_{1}=1$. We attribute this discrepancy to decoherence, which causes the qubit’s no-click trajectory to spiral into the Bloch sphere’s interior. This altered trajectory permits a forbidden region and a fixed point to exist in the presence of damped oscillations.
@@
At higher measurement strengths, a pronounced peak develops at the fixed point, consistent with the predicted divergence. By fitting $\tau_{\theta}(\theta)$ to the expected power law and extracting the zero crossing of $\xi$ (Fig. 4b), we locate the second transition at $\lambda^{\mathrm{obs}}_{2}=0.92\pm 0.01$ (Methods and Supplementary Information Section 7). Strikingly, this lies below the first transition, inverting the prediction of the ideal model. This counterintuitive result again reflects the influence of decoherence: analytical and numerical calculations show that decoherence shifts the second transition to lower $\lambda$ while raising the first, causing them to cross (Fig. 6 and Supplementary Information Section 6).
@@
Figure 4: Dwell time per unit angle $\tau_{\theta}(\theta)$ across the second transition. a, Ridge plot showing how the dwell time evolves with measurement strength $\lambda$. Purple dashed lines are fits to the analytical model $\tau_{\theta}(\theta)\sim\left[\theta-\theta_{+}(\lambda)\right]^{\xi(\lambda)}$ (Methods), performed only when $\theta_{+}>-\pi/2$. Inset: zoom-in on four representative values of $\lambda$ illustrating the transition: at $\lambda=0.4$, the dwell time is nonzero across all angles; at $\lambda=0.57$, a forbidden region has formed, before the first transition at $\lambda^{\mathrm{obs}}_{1}$; at $\lambda=0.88$, the dwell time remains small near $\theta_{+}$; at $\lambda=1.17$, a pronounced peak develops. Vertical arrows in both the main panel and the inset indicate the scale. b, Fitted critical exponent $\xi$ versus $\lambda$. Error bars represent $1\sigma$ confidence intervals. The dashed black line is a fit to the analytical model, yielding $\lambda^{\mathrm{obs}}_{2}=0.92\pm 0.01$ (vertical red line) at $\xi=0$ (horizontal grey line), marking the transition from finite to diverging dwell time near the forbidden region. Throughout, the coherent drive is $\Omega_{\mathrm{S}}/2\pi=100\,$kHz and the angular bin width is $2\pi/80$.
@@
Figure 5: Relaxation to the steady state across the third transition. At this transition, oscillatory decay gives way to overdamped relaxation; beyond it, stronger measurement slows relaxation, marking entry into the quantum Zeno regime. a, Ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$ for selected $\lambda$. Each data point is the average of $3000$ experimental runs. Error bars are smaller than the marker size. Dashed lines show fits using a sum of three complex exponentials. b, Extracted relaxation rates (blue) and oscillation frequencies (purple) versus $\lambda$. Error bars indicate $1\sigma$ confidence intervals. The dashed line is a fit of the oscillation frequency to the predicted behaviour (Methods). The transition is observed at $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$ (dashed red line), below the ideal value $\lambda^{\mathrm{c}}_{3}=2$ owing to finite waiting time in the detector’s excited state. Throughout, the coherent drive is $\Omega_{\mathrm{S}}/2\pi=100\,$kHz.
@@
The third and final transition occurs when the system enters the quantum Zeno regime [18, 38]. Beyond this transition point, measurement-induced backaction suppresses, rather than accelerates, the relaxation of the initial state towards the steady state [12, 16, 45, 29] (Figs. 2c,d). In addition, measurement-induced dephasing critically damps the ensemble dynamics at the transition point, turning decaying oscillations into an overdamped decay of the qubit polarisation. These changes stem from an exceptional point of the Liouvillian superoperator $\mathcal{L}$ in the master equation for the qubit’s density matrix $\hat{\rho}(t)$ [23, 9],

### v2

# Gradually opening Schrödinger’s box reveals a cascade of sharp dynamical transitions

@@ Abstract
Quantum mechanics predicts that unobserved systems may exist in a superposition of states, yet measurement produces definite outcomes—a tension at the heart of the quantum-to-classical boundary. How the transformation between these opposing regimes unfolds as observation strength increases has remained experimentally unexplored. Here, by continuously tuning the measurement strength on a superconducting qubit, we reveal that measurement-dominated dynamics emerge not gradually but through three distinct transitions: coherent oscillations abruptly halt; the system then freezes near a stable quantum state; and finally enters the quantum Zeno regime, where stronger observation paradoxically slows relaxation. Decoherence, rather than washing out this structure, reorganizes it—inverting the order in which transitions appear and decoupling signatures that coincide in idealized models. These results establish that the route from quantum dynamics to measurement-dominated behavior unfolds in sharp transitions governed by the interplay between observation and environment.
Measurement shapes the boundary between quantum and classical physics. Unobserved quantum systems evolve continuously under the Schrödinger equation, but measurement forces them into definite outcomes. Between these extremes lies a crossover whose internal structure — whether it is abrupt or gradual, featureless or rich — has remained experimentally unexplored. A continuously monitored two-level system offers a precisely controllable arena for resolving this question. Consider a qubit with ground state $\left|0\right\rangle$ and excited state $\left|1\right\rangle$, driven resonantly at a Rabi rate $\Omega_{\mathrm{S}}$ and monitored by a detector that clicks at a rate $\alpha$ when the system is in $\left|0\right\rangle$ (Fig. 1). A single dimensionless parameter, the measurement strength $\lambda\equiv\alpha/(2\Omega_{\mathrm{S}})$, captures the competition between coherent driving and the backaction of observation [34]. Quantum trajectory theory [5, 26, 45, 6] describes how stochastic detector clicks project the system to $\left|0\right\rangle$, while between clicks, the wavefunction evolves deterministically under an effective non-Hermitian Hamiltonian (supplementary text section 4[1, 3]
@@
The limiting regimes are well-understood. For weak measurements ($\lambda\ll 1$), coherent driving dominates, and the qubit undergoes Rabi oscillations (Fig. 1). For strong measurements ($\lambda\gg 1$), measurement backaction [16, 43, 28, 10] takes over: the sustained absence of clicks drives the qubit towards $\left|1\right\rangle$, producing a “continuous quantum jump” [22].
@@
Figure 2: Overview of dynamical transitions in the ideal monitored qubit model. As the dimensionless measurement strength $\lambda\equiv\alpha/(2\Omega_{\mathrm{S}})$ increases, the qubit dynamics evolve through four distinct regimes (A–D), separated by three sharp transitions at $\lambda^{c}_{1-3}=\{1,\frac{2}{\sqrt{3}},2\}$. Each row displays simulated dynamics in the different regimes, showing (from top to bottom) a representative binary click record, the conditional excited state probability $P_{\left|1\right\rangle}^{\mathrm{cond}}$ of the corresponding trajectory, and the ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$. The Bloch spheres show the average qubit dwell time per unit angle $\tau_{\theta}(\theta)$ (orange curves), plotted as radial height above the unit circle (a.u.), alongside the time evolution of the ensemble-averaged state initialized at $\left|0\right\rangle$ (teal curves). (A) For $\lambda<\lambda^{\mathrm{c}}_{1}$, trajectories exhibit periodic Rabi oscillations, intermittently reset to $\left|0\right\rangle$ by detector clicks, enabling the qubit to access the full range of polar angles. (B) The first transition at $\lambda^{\mathrm{c}}_{1}$ marks the onset of smooth, deterministic no-click evolution towards a stable eigenstate at polar angle $\theta_{+}$ (dotted gray lines), creating a forbidden region on the Bloch sphere (gray shading). Detector clicks typically interrupt this no-click evolution before the qubit reaches the eigenstate. (C) Beyond the second transition ($\lambda>\lambda^{\mathrm{c}}_{2}$), the qubit frequently reaches and freezes near the eigenstate before a click occurs, resulting in a divergent dwell time at $\theta_{+}$. (D) At the third transition point ($\lambda=\lambda^{\mathrm{c}}_{3}$), the ensemble-averaged dynamics shift from oscillatory behavior to overdamped decay with increasingly slow relaxation, marking the onset of quantum Zeno behavior. In realistic experimental conditions, decoherence modifies both the locations and the ordering of these transitions.
@@
A recent theoretical study [40] predicted that the crossover between these two regimes occurs through a cascade of sharp transitions, each unveiling qualitatively distinct dynamics and statistics of quantum trajectories (Fig. 2).
@@
Figure 3: Observation of the transition to the quantum jump regime. Below the first transition (bottom Bloch sphere), coherent driving (blue arrows) dominates measurement-induced backaction (red arrows) across the entire Bloch sphere, resulting in oscillations. Above it (top Bloch sphere), measurement backaction precisely balances the coherent drive at a specific angle, causing oscillations to halt and giving rise to quantum jump dynamics. (A) Tomographic data showing the excited state population $P_{\left|1\right\rangle}^{\mathrm{cond}}$ as a function of the no-click sequence duration and the measurement strength $\lambda$, revealing the shift from oscillatory to jump-like behavior. The transition is observed at $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$ (dashed red line). Each data point is the average of $1500$ experimental runs. (B) One-dimensional cuts below (bottom panel, $\lambda=0.5$) and above (top panel, $\lambda=2.2$) the first transition at $\lambda$ values indicated by the dotted lines in (A)). The gray lines show the simulated no-click dynamics with independently-measured parameters. Shot-noise error bars are smaller than the marker size. (C) Histograms of no-click durations further confirm this transition. We initialize the qubit in $\left|1\right\rangle$ (in contrast to $\left|0\right\rangle$ in A, B) to suppress the effect of errors in long no-click intervals. The gray solid lines show fits using a sum of three complex exponentials (Methods). Throughout this figure, the single-measurement integration time is $T_{\text{int}}=320\,$ns and the coherent drive strength is set at $\Omega_{\mathrm{S}}/2\pi=100\,$kHz. The tomographic measurement is preceded by a measurement to verify that the system is in the qubit manifold.
@@
Here, we observe this transition cascade experimentally using a superconducting circuit in which a system qubit is coupled to an ancillary detector qubit, forming a V-type three-level system similar to ref. [22]. We control the measurement strength $\lambda$ by varying the amplitude of a drive resonant with the detector qubit transition (see Methods). Tuning $\lambda$ reveals three dynamical transitions marking the emergence of measurement-dominated behavior (Fig. 2). At an exceptional point of the non-Hermitian Hamiltonian, coherent oscillations abruptly cease and trajectories begin evolving towards a stable state—the hallmark of continuous quantum jumps [22, 18]. Beyond a second threshold, the dwell time near this state diverges: trajectories “freeze” in its vicinity. A third transition signals entry into the quantum Zeno regime, where stronger measurement counterintuitively suppresses relaxation [32, 25, 12, 14].
@@
Our analysis combines binary detector click records with conditional quantum state tomography. This allows us to reconstruct the qubit’s dynamics conditioned on specific measurement outcomes, resolving dynamical features that ensemble averages would obscure. We find that decoherence does not merely shift the critical points predicted by the idealized theory, but fundamentally alters the transitions’ character and ordering. Nevertheless, the transitions themselves remain robust features of the dynamics. These results elucidate the emergence of continuous quantum jumps and provide a map of dynamical phases in a monitored qubit.
@@
The first transition, heralding the onset of continuous quantum jumps, arises from competition between two coherent processes acting during no-click evolution (equation (1)): the Rabi drive attempts to rotate the qubit around the Bloch sphere, while the no-click measurement backaction pulls it towards the excited state $\left|1\right\rangle$. Below a critical measurement strength $\lambda<\lambda^{\mathrm{c}}_{1}$, Rabi driving dominates at all polar angles $\theta$ on the Bloch sphere, allowing the qubit to complete full periodic rotations. As $\lambda$ approaches $\lambda^{\mathrm{c}}_{1}$, the oscillations become increasingly asymmetric, reflecting the growing impact of measurement backaction (Fig. 2A). The critical point $\lambda^{\mathrm{c}}_{1}=1$ represents the threshold at which measurement backaction first becomes strong enough to halt the qubit’s rotation. For any $\lambda\geq\lambda^{\mathrm{c}}_{1}$, backaction precisely balances coherent driving at an angle $\theta_{+}(\lambda)=2\arctan\!\left(\sqrt{\lambda^{2}-1}-\lambda\right)$, creating a stable fixed point [^1] A second fixed point appears at $\theta_{-}\equiv-\pi-\theta_{+}$ [40]. However, this fixed point is unstable, so the qubit reaches it only if specifically initialized in that state.  [35]. Trajectories now execute continuous, deterministic quantum jumps from $\left|0\right\rangle$ towards the eigenstate of $\hat{H}_{\text{eff}}$, $\left|\psi_{+}\right\rangle\equiv\sin\!\left(\frac{\theta_{+}}{2}\right)\left|0\right\rangle+\cos\!\left(\frac{\theta_{+}}{2}\right)\left|1\right\rangle$, until a stochastic detector click resets the qubit to $\left|0\right\rangle$. The fixed point at $\left|\psi_{+}\right\rangle$ creates a forbidden region $(-\pi,\theta_{+}(\lambda)]$ on the Bloch sphere, which eventually encompasses the entire left hemisphere as measurement strength increases [40] (see the Bloch spheres in Figs. 2B–D).
@@
This transition coincides with an exceptional point — a degeneracy in the spectrum of the non-Hermitian no-click Hamiltonian $\hat{H}_{\text{eff}}$ where eigenvalues coalesce. Since the Hamiltonian is parity-time ($\mathcal{PT}$) symmetric, the exceptional point separates unbroken and broken $\mathcal{PT}$-symmetric phases [3]. Beyond this point, the eigenvalue difference becomes purely imaginary, causing the oscillation frequency to vanish. Similar exceptional-point transitions have been observed in other systems [29, 46, 7, 8, 42].
@@
To observe this transition experimentally, we reconstruct the qubit’s evolution conditioned on the absence of detector clicks. We perform qubit tomography after detecting in real time a no-click sequence of the desired duration. Repeating this procedure across different no-click durations and measurement strengths reveals a clear shift from oscillatory to quantum jump dynamics (Figs. 3A,B). The measurements are in excellent agreement with numerical simulations of a model that incorporates decoherence (supplementary text section 8). We find that the Rabi oscillations become increasingly distorted as $\lambda$ approaches the critical value.
@@
The distribution of no-click durations independently confirms this transition. Initializing the system in $\left|1\right\rangle$, we record the waiting time to the first detector click. Since the instantaneous click rate varies as $r(\theta)\equiv\alpha\sin^{2}\!\left(\frac{\theta(t)}{2}\right)$, Rabi oscillations imprint a periodic modulation on this distribution. The data (Fig. 3C) confirm this expectation: oscillatory patterns at weak measurements give way to monotonic decay as $\lambda$ increases. From the vanishing of oscillations, we determine that the transition to quantum-jump dynamics occurs at $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$ (Methods and supplementary text section 7).
@@
A defining feature of strongly measured quantum systems is dynamical freezing: the qubit should not merely approach the eigenstate $\left|\psi_{+}\right\rangle$ but remain trapped in its vicinity for extended periods. To probe this behavior, we examine the dwell time per unit angle $\tau_{\theta}(\theta)$, defined such that $\tau_{\theta}(\theta)\Delta\theta$ gives the mean time spent in the angular interval $\theta,\theta+\Delta\theta)$. Near the fixed point at $\theta_{+}$, theory predicts a power-law scaling $\tau_{\theta}(\theta)\sim\left[\theta-\theta_{+}(\lambda)\right]^{\xi(\lambda)}$ [40], where the sign of the critical exponent $\xi$ reflects a competition between the click rate and the rate at which trajectories approach $\theta_{+}$. When clicks dominate ($\xi>0$), the dwell time vanishes at the fixed point (Fig. 2B). When the approach rate dominates ($\xi<0$), trajectories accumulate faster than clicks can reset them, and the dwell time diverges (Fig. 2C). This sign change defines a second critical threshold at $\lambda^{\mathrm{c}}_{2}=\frac{2}{\sqrt{3}}\approx 1.15$ in the ideal model (supplementary text section 4), marking the onset of dynamical state freezing.
@@
We reconstruct $\tau_{\theta}(\theta)$ experimentally by initializing the qubit in $\left|1\right\rangle$, recording the angular distribution $\rho(\theta)$ of first-click events, and normalizing by the instantaneous click rate $r(\theta)$ (Methods and supplementary text section 4). Fig. 4A shows the resulting dwell-time distributions across a range of measurement strengths. For weak measurement strengths ($\lambda\ll 1$), the qubit completes full Rabi oscillations and $\tau_{\theta}(\theta)$ is approximately uniform. As $\lambda$ increases, a forbidden angular region emerges. Intriguingly, this occurs at $\lambda\approx 0.57$, well before oscillations cease at $\lambda^{\mathrm{obs}}_{1}=0.99$. This decoupling of the two signatures of the first transition is absent in the ideal model, where both appear simultaneously at $\lambda^{\mathrm{c}}_{1}=1$. We attribute this discrepancy to decoherence, which causes the qubit’s no-click trajectory to spiral into the Bloch sphere’s interior. This altered trajectory permits a forbidden region and a fixed point to exist in the presence of damped oscillations.
@@
At higher measurement strengths, a pronounced peak develops at the fixed point, consistent with the predicted divergence. By fitting $\tau_{\theta}(\theta)$ to the expected power law and extracting the zero crossing of $\xi$ (Fig. 4B), we locate the second transition at $\lambda^{\mathrm{obs}}_{2}=0.92\pm 0.01$ (Methods and supplementary text section 7). Strikingly, this lies below the first transition, inverting the prediction of the ideal model. This counterintuitive result again reflects the influence of decoherence: analytical and numerical calculations show that decoherence shifts the second transition to lower $\lambda$ while raising the first, causing them to cross (Fig. 6 and supplementary text section 6).
@@
Figure 4: Dwell time per unit angle $\tau_{\theta}(\theta)$ across the second transition. (A) Ridge plot showing how the dwell time evolves with measurement strength $\lambda$. Purple dashed lines are fits to the analytical model $\tau_{\theta}(\theta)\sim\left[\theta-\theta_{+}(\lambda)\right]^{\xi(\lambda)}$ (Methods), performed only when $\theta_{+}>-\pi/2$. Inset: zoom-in on four representative values of $\lambda$ illustrating the transition: at $\lambda=0.4$, the dwell time is nonzero across all angles; at $\lambda=0.57$, a forbidden region has formed, before the first transition at $\lambda^{\mathrm{obs}}_{1}$; at $\lambda=0.88$, the dwell time remains small near $\theta_{+}$; at $\lambda=1.17$, a pronounced peak develops. Vertical arrows in both the main panel and the inset indicate the scale. (B) Fitted critical exponent $\xi$ versus $\lambda$. Error bars represent $1\sigma$ confidence intervals. The dashed black line is a fit to the analytical model, yielding $\lambda^{\mathrm{obs}}_{2}=0.92\pm 0.01$ (vertical red line) at $\xi=0$ (horizontal gray line), marking the transition from finite to diverging dwell time near the forbidden region. Throughout, the coherent drive is $\Omega_{\mathrm{S}}/2\pi=100\,$kHz and the angular bin width is $2\pi/80$.
@@
Figure 5: Relaxation to the steady state across the third transition. At this transition, oscillatory decay gives way to overdamped relaxation; beyond it, stronger measurement slows relaxation, marking entry into the quantum Zeno regime. (A) Ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$ for selected $\lambda$. Each data point is the average of $3000$ experimental runs. Error bars are smaller than the marker size. Dashed lines show fits using a sum of three complex exponentials. (B) Extracted relaxation rates (blue) and oscillation frequencies (purple) versus $\lambda$. Error bars indicate $1\sigma$ confidence intervals. The dashed line is a fit of the oscillation frequency to the predicted behavior (Methods). The transition is observed at $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$ (dashed red line), below the ideal value $\lambda^{\mathrm{c}}_{3}=2$ owing to finite waiting time in the detector’s excited state. Throughout, the coherent drive is $\Omega_{\mathrm{S}}/2\pi=100\,$kHz.
@@
The third and final transition occurs when the system enters the quantum Zeno regime [19, 40]. Beyond this transition point, measurement-induced backaction suppresses, rather than accelerates, the relaxation of the initial state towards the steady state [13, 17, 47, 30] (Figs. 2C,D). In addition, measurement-induced dephasing critically damps the ensemble dynamics at the transition point, turning decaying oscillations into an overdamped decay of the qubit polarization. These changes stem from an exceptional point of the Liouvillian superoperator $\mathcal{L}$ in the master equation for the qubit’s density matrix $\hat{\rho}(t)$ [24, 9],